Hardware for the HV system

The system is powered by 180 HV channels, 90 positive and 90 negative, by 15 A1534N boards and 15 A1534P boards, hosted in 2 SY1527 power supplies.

In March 2006 CAEN delivered the 2 Power Supplies and almost all the boards. These are the results of the tests performed on all boards.

CAEN Power Supplies
Power Supply name Rack Alias (proposal) Remarks
SY1527 alitofcae000 CR4-Z02 alitofsyhvn FIXME hosting negative boards
SY1527 alitofcae000 CR4-Z03 alitofsyhvp FIXME hosting positive boards
SY1527LC alitofcae000 I03 alitofsylv?? FIXME hosting 3 branch controllers
SY1527LC alitofcae000 I32 alitofsylv?? FIXME hosting 3 branch controllers
SY1527LC alitofcae000 O03 alitofsylv?? FIXME hosting 3 branch controllers
SY1527LC alitofcae000 O32 alitofsylv?? FIXME hosting 3 branch controllers

HV Positive Boards
HV Board Serial n. Test 2006/04/04-05 Test 2006/06/02 Remarks Remarks Nov.06
1 A1534P 20 not tested, used in lab 167 Ok prototype, refurbished by CAEN BAD .
2 A1534P 24 Ok Ok -- BAD (ch 0-1 switch off)
3 A1534P 25 Ok Ok -- --
4 A1534P 26 Ok, OvC /UnV in Ch02 once Ok -- BAD
5 A1534P 27 Ch01,Ch05 > 1A when OFF, Ch02 broken Ok repared by CAEN in May'06 --
6 A1534P 30 -- -- -- NEW: Under test
7 A1534P 31 Burned channels Ok repared by CAEN, May'06 Under test (again)
8 A1534P 32 Ok Ext-Disabled repaired by CAEN, 9-Aug'06 --
9 A1534P 33 HVMax not calibrated Ok HVMax checked by CAEN, May'06 --
10 A1534P 34 HVMax not calibrated Ok HVMax checked by CAEN, May'06 --
11 A1534P 35 HVMax not calibrated Ok HVMax checked by CAEN, May'06 --
12 A1534P 36 HVMax not calibrated Ok HVMax checked by CAEN, May'06 --
13 A1534P 37 Ch03 burned Ok repared by CAEN, May'06 --
14 A1534P 38 Ok Ok -- --
15 A1534P 39 Ok Ok -- --
16 A1534P 41 Ok, Ch00 UnV /OvC once Ok -- --
17 A1534P 42 -- -- -- NEW: Under test
18 A1534P 43 -- -- -- BAD
HV Negative Boards
HV Board Serial n. Test 2006/04/04-05 Test 2006/06/02 Remarks Remarks Nov.06
1 A1534N 21 possible spikes on Ch00 Ok -- --
2 A1534N 22 Ok Ok -- --
3 A1534N 23 -- -- -- NEW: Under test
4 A1534N 24 Ok Ok -- --
5 A1534N 25 -- -- -- NEW: Under test
6 A1534N 26 Ok Ok -- --
7 A1534N 27 Ok Ok -- --
8 A1534N 28 Ok Ok -- --
9 A1534N 29 Ok Ok -- --
10 A1534N 40 -- -- -- BAD
11 A1534N 41 -- -- -- NEW: Under test
12 A1534N 42 -- -- -- NEW: Under test
13 A1534N 43 HVMax not calibrated Ok -- --
14 A1534N 44 HVMax not calibrated Ok -- --
15 A1534N 45 HVMax not calibrated Ok -- --
16 A1534P 46 Ext-Dis Ok repared by CAEN, May'06 --
17 A1534N 47 Ok Ok -- --
18 A1534N 49 Ok -- used for cosmics since 2006/04/05 --
19 A1534N 50 HVMax not calibrated Ok -- --

15 KV Positive Boards
HV Board Serial n. Remarks
A1526P 49 Ok, 2006/04/04
A1526P 50 Ok, 2006/04/04
15 KV Negative Boards
HV Board Serial n. Remarks
A1526N 54 Ok, 2006/04/04
A1526N 56 Ok, 2006/04/04


20-Feb-2007

A1534N 21 and A1534P 32 given back to CAEN (S. Petrucci).

8-Mar-2007

Returned by Di Maio.


Test of HV distribution boxes

On 29/08/2006 (first test at 8 kV) and 31/08/2006 (second test at 12 kV) 15 distribution boxes have been tested.

Results:

  • 10 boxes OK (tested with 8 kV and 12 kV)
  • 5 boxes NOT_OK
    • 4 boxes failed the first test (OverCurrent around 5000-6000 V)
    • 1 box succeded the test at 8 kV but failed the test at 12 kV (and now it's failing also at 5000-6000 kV)

Test Procedure:

Testing the internal connections:

  1. GND connections: the 5 external shells of the unipolar cables and the central pin of the tripolar cable are connected
  2. Hot cable: the resistance between the single pin of the tripolar connector and the internal pin of the unipolar connectors is 1 MΩ
  3. the resistance between the internal pins of the unipolar connectors is 2 MΩ
  4. the "cold contact" should be connected to the external shield (:?: CHECK)

Test at 8000 V:

  • a tripolar cable conects the box to the A1534 board, and the channel is switched on with V0=8000 V
  • the box is kept in High Voltage for a few minutes.

Test at 12000 V:

  • a metallic box has being built to permit the connection from the A1526 board (unipolar connectors) to the red box (tripolar connectors)

Location:

  • all the material is now in room 10, blg 29.


2. Test of HV distribution boxes

On 14/09/2006 CPE delivered 9 boxes, without any painting.

Results:

  • 3 boxes OK at 12 kV (we tested one at 13 kV getting sparks, so we decided not to test the other two at 13).
  • 6 boxes NOT OK

The 6 have been dismantled to check if the origin of the sparks was visible, without success.

All the remaining boxes and the connection boxes (thanks to Dimitri for helping preparing them) are in Room 10.

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-06-25 - OmbrettaPinazza
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    AliceTOF All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback